# Towards Social AI: A Survey on Understanding Social Interactions

### v1

@@ V-B3 Facial Expression Datasets
V-C Datasets with Multimodal Cues
V-C1 Multimodal Emotion Datasets
As the field of multimodal emotion analysis progresses, a variety of sentiment analysis and emotion recognition datasets have been introduced to facilitate research and development in this area. All of these datasets consist of annotated data that offer video, audio, and transcribed texts as available modalities, while some of them also include labels for facial expressions or gestures. Classification categories vary across datasets; however, Ekman’s six universal emotions [301] (Joy, Sadness, Fear, Anger, Surprise, and Disgust) serve an important role in category design. Given the subjective and biased nature of emotion judgment, it is common practice to preset discrete emotion categories and merge annotation results from multiple annotators using a majority voting scheme. Additionally, Mean Absolute Error (MAE) and F1-Score are commonly used as evaluation metrics. In this section, we will provide a brief overview of four widely used multimodal datasets for sentiment/emotion analysis.
IEMOCAP [295] contains 12 hours of audio-visual data from 5 dyadic sessions acted by 10 actors. The data includes motion capture of facial expressions, dialog transcriptions, video, and audio recordings. The emotions are expanded from six universal emotions[301] to include fear, excitement, and surprise. Primitive attributes such as valence, activation, and dominance are also included.
SEMAINE The SEMAINE dataset [296] contains 959 conversations with Sensitive Artificial Listener characters. The data is audiovisual recordings of 150 participants with text transcriptions. Each clip is annotated by six to eight raters among 27 classes along 5 emotional dimensions.
MOSI [297] consists of 93 opinion videos from YouTube, featuring 89 distinct speakers (41 females and 48 males). The videos are segmented into 2,199 subjective and 1,503 objective opinion segments. Only the subjective segments are annotated with with sentiment annotations and intensity on a linear scale from +3 (strongly positive) to -3 (strongly negative) by two trained annotators.
MOSEI [298] is an extension of MOSI, with 23,453 video segments featuring 1,000 speakers and 250 topics. It adopts MOSI’s sentiment intensity labeling and adds emotion labels. Apart from incorporating the six Ekman emotions[301], it establishes a linear scale to indicate the possible presence of these emotions. Annotations are provided by three Amazon Mechanical Turk annotators.
MELD [215] is a multimodal extension of Emotionlines[42], sourcing data from the TV series “Friends”. It consists of 1,433 dialogues (13,000 utterances), including multi-party conversations. Ekman’s six emotions are expanded with Neutral and Non-Neutral labels. Additional sentiment classes (positive, negative, neutral) are utilized to further annotate complex emotions such as surprise.
V-C2 Conversation Dynamics Analysis Datasets
The dataset for conversation dynamics analysis may have a large variation since the dynamic nature of conversations can be observed in different aspects. Typically, we consider several datasets of social scenarios that include RGB frames and human’s speech. These datasets may vary in the tasks (such as classification, localization, etc.) but they are all used to study how the speech impacts human’s intention and attention as well as its dynamic nature.
SALSA [238] dataset contains 1-hour recordings of two unconstrained social scenes: a poster session and a cocktail party. Videos are captured from 4 synchronized viewpoints using a surveillance system, resulting in distant participants and noisy audio. The dataset includes audio, accelerometer data, and infrared proximity. Annotations consist of facing formulation, head and body orientation, and participant bounding boxes.
Triadic Interaction [241] is a 3D motion capture dataset containing 180 videos of a haggling game scenario with 120 participants. It captures triadic interactions in a lab setting, including video (500 views), depth (10 sensors), audio (23 mics), individual voice, and 3D point clouds. It provides benchmarks for predicting speaking status, social formation, and body gestures in multi-party interactions.
ConfLab The ConfLab [240] dataset is also established for F-formulation estimation similar to SALSA. The dataset comprises 1-hour overhead-view videos of people attending a conference. In addition to RGB videos, the dataset also collects audio signal, IMU data and Bluetooth proximity of via the wearable devices. The annotations include F-formulation, keypoints and speaking status of each participant.
Multi-speaker Conversation [235] contains egocentric video recordings of 50 participants divided into groups of 5, engaging in conversations around a table. Each video covers one group, and participants are instructed to listen or engage in their own group, allowing for the observation of attention labels. The dataset includes approximately 20 hours of egocentric videos with six-channel audio streams.
VideoGazeSpeech VideoGazeSpeech [233] is a recent dataset that investigates the correlation of speech and human’s gaze in social scenarios. The dataset is built based on the data collected by Xu et al. [302], but VideoGazeSpeech only selects 29 videos with audio stream. Each video has 400 500 frames involving 2 4 people on average. Frame-level gaze target is provided on each frame.
Werewolf Among Us (Exp.) [232] is an expanded dataset of [13]. This dataset provides new benchmarks and labels for speaking target identification, pronoun coreference resolution, and mentioned player prediction. It enables a high level of understanding of multimodal conversation (e.g., verbal, non-verbal cues) in multi-party settings. It provides labels for the 234 videos in the original Werewolf Among Us dataset.
V-C3 Social Situation Analysis Datasets
There are social situation analysis datasets mainly for social relationship reasoning and social question answering. The performance of a model for social relation inference is commonly evaluated by classification accuracy. In terms of social question answering, a question is raised based on the social situation and a couple of answer candidates are provided. The AI models have to correctly parse the social scenes and select the correct answer(s) for the given question. The performance is usually measured by accuracy, i.e., the percentage of questions that are correctly answered. A variant for the question answering setting is generating answers in free-form language to the given questions instead of selecting from pre-defined answers. The performance is measured by BLEU [303], METEOR [304], ROUGE [305], BERTscore [306] and user study. Another dataset using werewolf games to evaluate the performance of models in learning the human’s persuasion strategies and voting intentions [13]. They straightforwardly use accuracy as the metric for the two tasks.
MovieGraphs [299] contains 7637 movie clips annotated with graphs that capture who is present, their attributes, interactions, and relationships. Each graph consists of several node types including characters, their attributes, relationships, interactions, topics of interactions, reasons behind interactions, and time stamps. The dataset includes free-text labels and encompasses about 100 relationships, such as parent, spouse, and friend.
TVQA [258] is a dataset for video question answering that is built on 6 TV shows spanning 3 genres: crime shows, sitcoms and medical dramas. The dataset consists of 152.5K QA pairs that span 21,783 video clips from 925 episodes. Each video has 7 questions with 5 answer candidates, and only one is the correct answer. The length of the videos clips vary in between 60-90 seconds containing social interaction and activities.
ViSR The Video-based Social Relation dataset (ViSR) [249] contains over 8,000 video clips annotated with 8 social relation types derived from social domain theory. The dataset covers a variety of daily life social relations such as parent-offspring, couple, friends, colleagues etc. The videos are from movies and TV shows, providing diverse scenes and backgrounds.
Social-IQ [242] consists of a set of 7500 questions, 52500 answers (30000 correct and 22500 incorrect) covering a broad range of 1250 social videos. The videos in this dataset come from a diverse set of YouTube videos consisting of distinct characters across the videos. The questions and answers in this dataset are related to multiple modalities in social events. This dataset was further extended to Social-IQ 2.0 [307].
LVU [300] evaluates understanding of long-form content through 9 tasks on 30,000 video clips up to 3 minutes. Tasks include content analysis (e.g., relationship, speaking style), user engagement prediction (e.g., ratio, popularity), and movie metadata prediction (e.g., genre, director). LVU contains 4 relationship types: ’friends’, ’wife-and-husband’, ’boyfriend-and-girlfriend’, and ’ex-boyfriend-and-ex-girlfriend’.
SMILE The SMILE [259] dataset is a multimodal dataset for understanding laughter in videos with language models. The dataset comprises 887 video clips, each paired with a language description about the reason for laughter for the corresponding video clip. The video clips come from TED talks, where each video is between 10-90 seconds long, and sitcoms, where each video is 7-60 seconds long.
Werewolf Among Us [13] contains 252 videos and transcripts of social deduction games. One Night Ultimate Werewolf and The Resistance: Avalon games are obtained from YouTube and Ego4D [146]. Players form two teams and use persuasion strategies. Six typical persuasion strategies are pre-defined on the dataset. Annotations include persuasion strategies per utterance, player roles (start and end), and voting outcomes.
VI Challenges and Future Directions
@@
VI-A Understanding Multimodal Social Cues
VI-A1 Multimodal Alignment
@@ VI-A2 Multimodal Integration
VI-A3 Multimodal Long-range Dependencies
@@
Social interactions often unfold over extended periods, with long-range dependencies between multimodal cues. For example, a provocative gesture made early in a conversation can affect the overall mood and influence subsequent verbal exchanges. However, most current multimodal social reasoning tasks rely on short videos or a few utterances [235, 231, 232, 242]. There are works that deal with long-term understanding in social contexts [300]. Still, their target video length is a few minutes. Truly effective social contextualization may require a comprehensive understanding of much longer lengths. Future benchmarks and techniques should be designed to address the long-term context in multimodal interaction data. Tasks like reasoning about outcomes in social deduction games [13], which involve lengthy conversations and complex dynamics, may serve as effective testbeds. Potential technical directions include memory augmentation [48, 313, 314] to efficiently capture long-range dependencies.
VI-A4 Multimodal Understanding Benchmarks.
@@
While several datasets have been introduced for multimodal social understanding, such as Social-IQ [242], Ego4D [146], MovieGraphs [299], and Werewolf-Among-Us [13, 232], there remains a need for more comprehensive benchmarks that fully capture the complexity of multimodal interactions. Existing datasets, while valuable, often fall short in certain aspects. For instance, Social-IQ provides a diverse range of social scenarios. However, there can be some biases between questions and answers [260] due to Q&A with multiple choice, which allows machines to sometimes solve problems without leveraging multiple modalities. Ego4D offers rich egocentric data but is limited in addressing the camera-wearer’s own multimodal cues (e.g., their face cannot be observed). MovieGraphs provides detailed graph-based annotations of social situations but the scenes are mainly staged and may not fully represent real-world, natural interactions. The Werewolf-Among-Us dataset captures complex group dynamics including multimodal social cues but are confined to specific game contexts. Future efforts should focus on creating large-scale datasets with diverse social interactions and fine-grained multimodal annotations across different modalities (e.g., text, audio, video). Moreover, the datasets should include challenging reasoning tasks that require the fine-grained multimodal alignment and integration of multimodal information over extended periods.
VI-B Understanding Multi-party Dynamics
VI-B1 One-to-many Dynamics
@@
Real social scenarios often involve individuals engaging in one-to-many interactions, such as a speaker addressing multiple listeners simultaneously. However, existing works mainly focus on one-to-one interaction dynamics [236, 232] even in multi-party environments. Modeling one-to-many dynamics is crucial for capturing the full complexity of social interactions. Future research should develop techniques to represent and reason about the relationships and dependencies between an individual and multiple interaction partners. This may involve creating new benchmarks specifically designed for one-to-many scenarios or extending graph neural networks [236, 231] to handle multiple interaction targets.
VI-B2 Group-level Dynamics
@@
VI-B3 Interaction History Modeling
@@
Current conversation dynamics is influenced by their interaction history between participants. In multimodal contexts, this interaction history encompasses not only verbal exchanges but also non-verbal cues such as gaze, facial expressions, and gestures. For instance, knowing who spoke to whom, who looked at whom, and who reacted to whom in previous interactions can provide important clues about ongoing conversation dynamics and interpersonal relationships. Future research can explore methods for encoding and integrating multimodal interaction history into multi-party models. This may involve extending temporal graph embeddings [317, 52] to incorporate multiple modalities in encoding interaction history.
VI-B4 Multi-party Understanding Benchmarks
@@
While some datasets have been introduced for multi-party interaction analysis such as Social-IQ [242] and Werewolf-Among-Us [13, 232], there are still gaps in the comprehensiveness and diversity of these benchmarks for studying multi-party dynamics, which is corroborated by findings from [18]. The Social-IQ dataset, although containing multi-party social videos, primarily focuses on question-answering tasks and lacks fine-grained annotations for modeling the relationships and interactions between participants. The Werewolf-Among-Us datasets, while capturing complex and fine-grained multi-party dynamics, is limited to a specific game scenario and may not generalize to other social contexts. An ideal multi-party interaction benchmark should cover a wide range of real-world multi-party scenarios, such as group discussions, team meetings, and social gatherings, with varying numbers of participants and group structures. The dataset should include rich annotations for tasks like identifying the referents in utterances, recognizing the relationships between participants, and predicting the group dynamics such as alliances or conflicts. Moreover, the benchmark should include longitudinal interaction data to facilitate the study of interaction history and its impact on group dynamics over time. This would allow researchers to investigate how relationships and roles evolve and how past interactions influence current behavior.
VI-C Understanding Beliefs
VI-C1 Cultural Awareness
@@
Cultural background significantly shapes beliefs, norms, and expectations, influencing their behavior and interpretation of social interactions. However, current multimodal interaction understanding approaches often overlook cultural aspects. Future research should prioritize cultural awareness in social interaction models by incorporating cultural knowledge and adapting to cultural differences. Future research may extend techniques like culture-aware language inference [318] in multimodal aspects including non-verbal social cues.
VI-C2 Relationship Awareness
@@
The relationship between interacting individuals, such as their social roles, power dynamics, or emotional connections, affects their shared beliefs and plays a significant role in shaping interaction dynamics. However, current multimodal approaches often treat all interactions uniformly without considering underlying relationships. Even though there exist works that deals with social relationship inference [250, 251, 252], most of them do not effectively incorporate predicted social relationships to interpret further social interactions. Future work should develop techniques integrating relationship knowledge between interacting parties in social reasoning. This may involve extending methods like relational graph neural networks [319, 248] that can learn relationship-aware representations. Moreover, future research also can explore ways to incorporate external knowledge about social roles and relationships into multimodal interaction models.
VI-C3 Individual Belief Awareness
@@
Understanding an individual’s beliefs, intentions, and mental states is crucial for interpreting their actions and predicting their behavior in social interactions. Current methods for individual belief modeling primarily rely on language-based cues. Future research should focus on developing techniques that integrate both verbal and non-verbal dynamics to infer individual beliefs and mental states from multimodal interaction data. This may involve approaches like theory of mind modeling [320], which aims to simulate cognitive processes of perspective-taking and mental state attribution, or inverse reinforcement learning [321] methods that can infer an individual’s goals and preferences from their observed behavior.
VI-C4 Belief-aware Benchmarks
@@
Existing multimodal interaction datasets often lack the diversity and granularity of annotations needed to effectively train and evaluate models that incorporate individual beliefs, social relationships, and cultural contexts. For example, while the Triadic Belief dataset [294] provides some annotations related to beliefs, it is limited to a narrow set of belief types in a specific lab-based interaction scenario and does not capture the full range and complexity of beliefs that shape real-world social interactions. Similarly, datasets like Social-IQ [242] and Werewolf-Among-Us [13, 232], although valuable for studying social interactions, do not provide detailed annotations about the beliefs, intentions, and mental states of individual participants, nor do they explicitly capture information about social relationships or cultural backgrounds that influence those beliefs. An ideal belief-aware benchmark should encompass a diverse range of real-world social interaction scenarios with varying cultural contexts, social relationships, and individual characteristics. Crucially, it should include rich, fine-grained annotations for various aspects of beliefs at both the individual and collective levels. This would include labels for intentions, goals, and mental states of each participant throughout the interaction, as well as annotations for the relationships between participants and their cultural backgrounds.
VII Conclusion
@@
In this survey, we provide a comprehensive overview of the current state of research in understanding social interactions through artificial intelligence. We examine the landscape of existing work across three key modalities: verbal cues, non-verbal cues, and multimodal approaches. Our analysis highlights the importance of three core capabilities for effective social understanding: multimodal integration, multi-party modeling, and belief awareness. We review relevant datasets and benchmarks, identifying their strengths and limitations. Looking ahead, we outline several critical challenges and promising future directions in each of these areas. These include improving multimodal alignment and integration, modeling complex group dynamics, and incorporating cultural and individual belief awareness into social interaction models. As the field progresses, addressing these challenges will be crucial for developing AI systems capable of truly understanding and seamlessly participating in human social interactions. By advancing along these research directions, we can move closer to the goal of creating socially intelligent AI that can effectively interpret, reason about, and engage in the multifaceted nature of human social interactions.

### v2

@@ 4.3.5 Discussion
4.4 Datasets with Multimodal Cues
Table III shows a comprehensive overview, comparing the attributes of datasets with multimodal cues. Further details for each dataset are included in the supplementary material.
5 Challenges and Future Directions
@@
5.1 Understanding Multimodal Social Cues
5.1.1 Multimodal Alignment
@@ 5.1.2 Multimodal Integration
5.1.3 Multimodal Long-range Dependencies
@@
Social interactions often unfold over extended periods, with long-range dependencies between multimodal cues. For example, a provocative gesture made early in a conversation can affect the overall mood and influence subsequent verbal exchanges. However, most current multimodal social reasoning tasks rely on short videos or a few utterances [251, 247, 248, 258]. There are works that deal with long-term understanding in social contexts [282]. Still, their target video length is a few minutes. Future research should be designed to address the long-term context in multimodal interaction data. Tasks like reasoning about outcomes in social deduction games [11], which involve lengthy conversations and dynamics, may serve as effective testbeds. Potential directions include memory augmentation [45, 288, 289] to efficiently capture long-range dependencies.
5.1.4 Multimodal Understanding Benchmarks.
@@
While several datasets have been introduced for multimodal social understanding, there remains a need for more comprehensive benchmarks that fully capture the complexity of multimodal interactions. For instance, Social-IQ [258] provides a diverse range of social scenarios. However, there can be some biases between questions and answers [276] due to Q&A with multiple choice, which allows machines to sometimes solve problems without leveraging multiple modalities. Ego4D [147] offers rich egocentric data but is limited in addressing the camera-wearer’s own multimodal cues (e.g., their face cannot be observed). MovieGraphs [281] provides detailed graph-based annotations of social situations but the scenes are mainly staged and may not fully represent real-world, natural interactions. Werewolf-Among-Us [11, 248] captures complex group dynamics including multimodal social cues but are confined to specific game contexts. Future efforts should focus on creating large-scale datasets with diverse social interactions and fine-grained multimodal annotations across different modalities (e.g., text, audio, video). Moreover, the datasets should include challenging reasoning tasks that require the fine-grained multimodal alignment and integration of multimodal information over extended periods.
5.2 Understanding Multi-party Dynamics
5.2.1 One-to-many Dynamics
@@
Real social scenarios often involve individuals engaging in one-to-many interactions, such as a speaker addressing multiple listeners simultaneously. However, existing works mainly focus on one-to-one interaction dynamics [252, 248] even in multi-party environments. Modeling one-to-many dynamics is crucial for capturing the full complexity of social interactions. Future research should develop techniques to represent and reason about the dependencies between an individual and multiple interaction partners. This may involve creating new benchmarks specifically designed for one-to-many scenarios or extending graph neural networks [252, 247] to handle multiple interaction targets.
5.2.2 Group-level Dynamics
@@
5.2.3 Interaction History Modeling
@@
Current conversation dynamics is influenced by their interaction history between participants. For instance, knowing who spoke to whom, who looked at whom, and who reacted to whom in previous interactions can provide important clues about ongoing conversation dynamics and interpersonal relationships. Future research can explore methods for encoding and integrating multimodal interaction history into multi-party models. This may involve extending temporal graph embeddings [292, 48] to incorporate multiple modalities in encoding interaction history.
5.2.4 Multi-party Understanding Benchmarks
@@
While some datasets have been introduced for multi-party interaction analysis, there are still gaps in the comprehensiveness and diversity of these benchmarks, which is corroborated by findings from [16]. The Social-IQ dataset [258], although containing multi-party social videos, primarily focuses on question-answering tasks and lacks fine-grained annotations for modeling the relationships and interactions between participants. Werewolf-Among-Us dataset [11, 248], while capturing complex and fine-grained multi-party dynamics, is limited to specific game scenarios and may not generalize to other social contexts. An ideal multi-party interaction benchmark should cover a wide range of real-world multi-party scenarios, such as group discussions, team meetings, and social gatherings, with varying numbers of participants and group structures. The dataset should include rich annotations for tasks like identifying the referents in utterances, recognizing the relationships between participants, and predicting the group dynamics such as alliances or conflicts. Moreover, the benchmark should include longitudinal interaction data to facilitate the study of interaction history and group dynamics over time.
5.3 Understanding Beliefs
5.3.1 Cultural Awareness
@@
Cultural background significantly shapes beliefs, norms, and expectations, influencing their behavior and social interactions. However, current multimodal interaction understanding approaches often overlook cultural aspects. Future research should prioritize cultural awareness in social interaction models by incorporating cultural knowledge and adapting to cultural differences. Future research may extend techniques like culture-aware language inference [293] in multimodal aspects including non-verbal social cues.
5.3.2 Relationship Awareness
@@
The relationship between interacting individuals, such as their social roles, power dynamics, or emotional connections, affects their shared beliefs and plays a significant role in shaping interaction dynamics. However, current multimodal approaches often treat all interactions uniformly without considering underlying relationships. Even though there exist works that deals with social relationship inference [266, 267, 268], most of them do not effectively incorporate predicted social relationships to interpret further social interactions. Future work should develop techniques integrating relationship knowledge between interacting parties in social reasoning. This may involve extending methods like relational graph neural networks [294, 264] that can learn relationship-aware representations. Moreover, future research also can explore ways to incorporate external knowledge about relationships into social understanding.
5.3.3 Individual Belief Awareness
@@
Understanding an individual’s beliefs, intentions, and mental states is crucial for interpreting their behaviors in social interactions. Current methods for individual belief modeling primarily rely on language-based cues. Future research should focus on developing techniques that integrate both verbal and non-verbal dynamics to infer individual beliefs and mental states from multimodal interaction data. This may involve approaches like theory of mind modeling [295], which aims to simulate cognitive processes of perspective-taking and mental state attribution, or inverse reinforcement learning [296] methods that can infer an individual’s goals and preferences from their observed behavior.
5.3.4 Belief-aware Benchmarks
@@
Existing multimodal interaction datasets often lack the diversity and granularity of annotations needed to effectively train and evaluate models that incorporate individual beliefs, social relationships, and cultural contexts. For example, while the Triadic Belief dataset [297] provides some annotations related to beliefs, it is limited to a narrow set of belief types in specific lab-based interaction scenarios. Similarly, datasets like Social-IQ [258] and Werewolf-Among-Us [11, 248], although valuable for studying social interactions, do not provide detailed annotations about the beliefs, intentions, and mental states of individual participants, nor do they explicitly capture information about social relationships or cultural backgrounds that influence those beliefs. Ideal belief-aware benchmarks should encompass a diverse range of real-world social interaction scenarios with varying cultural contexts, social relationships, and individual characteristics. Crucially, it should include fine-grained annotations for various aspects of beliefs at both the individual and group levels. This would include labels for intentions, goals, and mental states of each participant throughout the interaction, as well as annotations for the relationships between participants and their cultural backgrounds.
6 Conclusion
@@
In this survey, we provide a comprehensive overview of the current state of research in understanding social interactions through artificial intelligence. We examine the landscape of existing work across three key modalities: verbal cues, non-verbal cues, and multimodal approaches. Our analysis highlights the importance of three core capabilities for effective social understanding: multimodal integration, multi-party modeling, and belief awareness. Looking ahead, we outline several critical challenges and promising future directions in each of these areas. These include improving multimodal alignment and integration, modeling complex group dynamics, and incorporating cultural and individual belief awareness into social interaction models. As the field progresses, addressing these challenges will be crucial for developing AI systems capable of truly understanding and seamlessly participating in human social interactions. By advancing along these research directions, we can move closer to the goal of creating socially intelligent AI that can effectively interpret, reason about, and engage in the multifaceted nature of human social interactions.
Acknowledgments
This work was supported in part by NSF CNS-2308994.